Stellar Corporation had 138,000 shares of stock outstanding on January 1, 2017. On May 1, 2017, Stellar issued 60,000 shares. On July 1, Stellar purchased 10,080 treasury shares, which were reissued on October 1.

Compute Stellar’s weighted-average number of shares outstanding for 2017.
Answer #1

Calculation of weighted average number of shares:

1) 138000*4/12= 46000

2) (138000-60000)*8/12= 26000

3) 10080*3/12= 2520

Weighted number of shares= 46000+26000+2520= 74520